OpenStudy (anonymous):

Find dy/dx by implicit differentiation. cot(y) = 3x − 6y

OpenStudy (anonymous):

can someone please help me with this?

OpenStudy (anonymous):

lets see if i still got my skills *knuckle crack*

OpenStudy (anonymous):

haha alright.

OpenStudy (anonymous):

does anything come to mind:)

OpenStudy (anonymous):

differentiate both sides and since y is a function of x we get: (-csc^2(y))y'=3-6y' now we get all the y's on one side... 7y'=3/(-csc^2(y)) y'= -3/7csc^2(y)

OpenStudy (anonymous):

pretty sure that's how that one goes

OpenStudy (anonymous):

pretty much just treat y like it's a general function instead of a variable and see what happens. lol.

OpenStudy (anonymous):

y'= -3/7csc^2(y)...so thats the answer..we learnt it in lecture but i was havng trouble with it

OpenStudy (anonymous):

yep. that's what i get anyway. does the back of the book give you the answer? if you want i can try to help you out with it more

OpenStudy (anonymous):

thats the wrong answer

OpenStudy (anonymous):

see our homework is submitted online so it marked it wrong but i have another shot

OpenStudy (anonymous):

then i fail. lol. my skills have faded

OpenStudy (anonymous):

lets see if i can figure out what i did wrong, hold on a sec

OpenStudy (anonymous):

haha its oki

OpenStudy (anonymous):

alright thanx:)

OpenStudy (anonymous):

ok, i know why i fail at life now. i differentiated correctly, but solved for y' incorrectly. super algebra fail. lets try this again: -csc^2(y)(y')=3-6y' y'[(-csc^2(y)-6)]=3 y'=3/(-csc^2(y)-6) or y'=-3/(csc^2(y)+6)